Grade 11Inorganic Chemistry

The wavelength of third line of balmer series for hydrogen atom is
A-21/100Rh
B-100/21Rh
C-21Rh/100
D-100Rh/21

Santosh Kumar
6 Years agoGrade 11

1 Answer

Vikas TU
6 Years ago
Dear student 
1/ lambda = Rh ( 1/22 – 1/52)
1/lambda = Rh (1/4 – 1/25)
1/lambda = 21/100  Rh 
lambda = 10Rh/21